OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0).

OpenStudy (pfenn1):

\[y=mx+b\]where m=slope and b=y-intercept \[m=\frac{y _{2}-y _{1}}{x _{2}-x _{1}}\]

OpenStudy (pfenn1):

Once you get m, choose one of the points, substitute the values for m, x, and y into the equation, and solve for b.

OpenStudy (anonymous):

y − 4 = −1/2(x + 6) y − 4 = 2(x + 6) y + 6 = −1/2(x − 4) y + 6 = 2(x − 4) can you choose one of these

OpenStudy (pfenn1):

Oops. I gave you the slope-intercept form of the equation, not the point-slope form. The point slope form is given by\[y – y_1 = m(x – x_1)\] You would still use the equation I gave you above for m. And then pick one of the points and substitute into the equation.

OpenStudy (turingtest):

@SOUTHERNPRIDE hi and welcome to Open Study I hope we can help you understand math better, but please don't request pure answers with no explanation; that is not the aim of this site. pfenn1 has given you enough information to solve the problem, can you tell us where you are having trouble?

OpenStudy (anonymous):

does it go like this? (-6,4)=(x1,y1) (2,0)= (x,y)

OpenStudy (pfenn1):

Close.\[(-6,4)=(x_1,y_1) ; (2,0)=(x_2,y_2) \]

OpenStudy (anonymous):

thank you that really helped

OpenStudy (pfenn1):

you're welcome